Determine the surface area generated by revolving the curve y=1/3 x³ over the interval (0,2), about the x-axis.


1
Expert's answer
2021-05-07T08:13:15-0400

Solution

If the curve y = f(x), a≤x≤b is rotated about the x-axis, then the surface area is given by

"A=2\pi\int_a^bf(x)\sqrt{1+[f'(x)]^2}dx"

In this case f(x) = 1/3 x³, f’(x) = x2 ,a = 0, b = 2 

"A=2\pi\int_0^2\frac{1}{3}x^3\sqrt{1+x^4}dx"

Substitution y = x4 

"A=\frac{\pi}{6}\int_0^{16}\sqrt{1+y}dy=\frac{\pi}{9}(1+y)^{3/2}|_0^{16}=\frac{\pi}{9}(17^{3/2}-1)=24.118"

Answer

"A=\frac{\pi}{9}(17^{3/2}-1)=24.118"